# PoI: Pixel of Interest for Novel View Synthesis Assisted Scene Coordinate Regression

### v2

@@ 3.3 Filtering strategy
We implement filtering operations at the intermediate stage of training. The filter will remove novel features of outlier prediction. The remaining features are the so-called FoI, and the corresponding pixels of FoI are the so-called PoI. Figure 3 shows an example of the PoI results in the 7Scenes and Cambridge Landmarks datasets. After filtering, we combine and shuffle the features $F$ and FoI and put them into the scene-specific MLP Head to estimate the scene coordinates.
@@
It is worth mentioning that we have set a dynamic weight for the loss of rendering pixels. Because in the early stage of training, we want the model to converge quickly. After determining the PoI, we gradually reduce the weight of the loss of PoI from 1 to 0.01, while for the pixels from query images, we set the weight to 1 during the entire training process.
@@
where T denotes traing data, $\tilde{\omega}$ denotes the dynamic weight of PoI loss changing from $\omega_{max}$ (set 1) to $\omega_{min}$ (set 0.01). $I_{iter}$ denotes the current iteration number and $N_{iter}$ is the total iterations. All rendering data is initially set as PoI. As the training progresses, we rule out outlier prediction points from PoI. At the end of the training, the choice of PoI and the loss weight of PoI are fixed.
In PoI, we would sample novel camera poses and render the corresponding images according to the images and the corresponding camera poses from the training set. In existing novel view synthesis supported visual localization, we usually have to balance the novel poses’ diversity and the images’ overall rendering quality. However, we do not need an overall well-rendered image in the PoI task because of the pixel-level optimization and filtering algorithm. Therefore, we should try to expand the diversity of novel poses. We use a unified sampling method for camera pose translation: grid sampling. The boundaries of the grid are calculated based on the camera pose of the training data; that is, the maximum and minimum values of the grid (x, y, z) are determined by the maximum and minimum values of the translations of all camera poses. We add a random perturbation to the rotation. The original rotation of each grid starts from the closest camera pose to that grid in the training data.
Figure 3: An example of the results of PoI in dataset 7Scenes and Cambridge Landmarks. To highlight the determined pixels of interest, we scale up the ‘Value’ (V) of the HSV representation of the images.
3.4 Applicable conditions and limitations
For the end-to-end SCR approach, PoI is difficult to use as a plug-and-play module, which is the limitation of the PoI method. First, end-to-end SCR methods use image-level loss as a supervision and have no pixel-level performance, which makes PoI unusable in these methods. Furthermore, even for two-stage SCR methods (init+e2e) such as DSAC*, all pixels of the same image should be supervised within one iteration in the e2e stage. If we filter out some pixels of the rendered images, aligning the rendered features and designing a differentiable RANSAC algorithm is difficult. Finally, pixel-wise shuffling (which is difficult to achieve in e2e) is also an important factor; without shuffling, poorly rendered pixels are more likely to appear in a batch of data. As a result, the network is more likely to get stuck at a local minimum.
For non-end-to-end training methods like ACE and GLACE, the difference is that we can easily shuffle pixels from all rendered images and query images because the supervision relies only on the camera intrinsics and the planar coordinates of each pixel without further requirements of per-frame joint supervision.
Take GLACE as an example; our PoI could also be used as in Figure 2.(b); the difference is that the backbone should be replaced. We use the global encoder of GLACE to extract the same dimension of global features as the ACE features. We add the global feature and the ACE feature together and get the target feature maps. The following procedure remains unchanged.
3.5 Sparse input
Sparse input visual localization is a challenging task since both CPR and SCR are not good at estimating unseen parts of the scene because the regression models are trained only from RGB with weak geometric constraints. However, with the help of sparse input NVS, we would be able to extend the training views.
In this case, the challenge is that rendered frames from sparse-view NVS are generally of lower quality compared with those from dense-input NVS. The ratio between the input views and the NVS views is also reduced. If we still use the raw PoI method, the implicit neural map will be mainly contributed by rendered pixels. The accuracy will be influenced. To address this problem, we propose a coarse-to-fine training approach as shown in Figure 4. In the coarse stage, we use the same setting as in PoI; the only difference is that all training data are rendered images. So, the filter is applied to all rendered pixels, which we call the self-pruning step. In this step, in order to retain adequate pixels for training, we raise the filter’s threshold (for reprojection errors). We get a coarse model after self-pruning training. In the refinement stage, we fine-tuned the mapping model using real data and the remaining rendered data; we set the learning rate to lower than that of the coarse stage throughout the fine-tuning process. In this step, all pixels are put into the model without filtering.
We finally get the fine-tuned model and experiment on both indoor and outdoor datasets. The results and implementation details can be found in chapter 4.3.
@@ 3.5 Sparse input
4 Experiment
Table 1: Median errors of camera pose regression methods and scene coordinate regression methods on the 7Scenes dataset Shotton et al. (2013). We bold the best results.
@@ 4 Experiment
4.1 Implementation details
Table 2: Results on Cambridge Landmarks, because of the obvious gap between SCR-based methods and CPR-based methods, we only list SCR-based methods. column ‘Mapping time’ shows the training time of these methods (the GS model can be trained offline, and the training time of the GS model is less than PoI), and column ‘Mapping size’ is the memory consumption for saving the parameters of the network. We bold the best result for group ‘SCR’ and group ‘SCR w/ glob’ separately.
@@ 4.1 Implementation details
SCR SANet Yang et al. (2019) Yes  1min  260M 32/0.5 32/0.5 10/0.5 16/0.6 328/2 84/0.8
SRC Dong et al. (2022) Yes 2min 40M 39/0.7 38/0.5 19/1 31/1.0 81/0.5 42/0.7
DSAC* Brachmann and Rother (2021) No 15h 28M 18/0.3 21/0.4 5/0.3 15/0.6 34/0.2 19/0.4
Poker Brachmann et al. (2023) No 20min 16M 18/0.3 25/0.5 5/0.3 9/0.3 28/0.1 17/0.3
EGFS Liu et al. (2025) No 21min 9M 14/0.3 28/0.1 19/0.4 5/0.2 10/0.3 15/0.3
GLACE Wang et al. (2024) No 20min 13M 19/0.3 17/0.4 4/0.2 9/0.3 19/0.1 14/0.3
NRP LENS Moreau et al. (2022b) No - - 33/0.5 44/0.9 27/1.6 53/1.6 - -
GSplatLoc Sidorov et al. (2024) No - - 27/0.46 20/0.71 5/0.36 16/0.61 - -
@@
Dataset We evaluate the performance of our approaches on two public datasets, Microsoft 7Scenes Shotton et al. (2013) (with SFM Pseudo GT) and Cambridge Landmarks Kendall et al. (2015). Our network takes RGB images and the camera extrinsic as input without using the depth information from the 7Scenes dataset or the reconstruction information from the Cambridge Landmarks dataset. We take the original resolution for the RGB images to make an accurate pose estimation.
To save time and computing resources, we do pose sampling and synthesis of new views offline and save the sampled camera poses and the rendered images on disk. During training time, we load rendered images along with the training set from the disk. For the scene ’kitchen’ only, we split the training data into two clusters using the camera poses. We follow the rule of poker ( a variant of ACE) and train two models with the clusters. During the evaluation, we choose the estimated pose from the model with a more significant number of inlier pixels of the Ransac algorithm. We use 1 NVIDIA V100 GPU for PoI training; for GLPoI, we use 4 V100 with distributed data-parallel training.
4.2 Localization results
The comparison of median translation and rotation errors between our proposed methods with different camera pose regression methods (top), scene coordinate regression methods (middle), and neural render pose estimation methods (bottom) in dataset 7Scenes is shown in Table 1. Generally speaking, scene coordinate regression methods outperform absolute camera pose regression methods in both translations and orientations. Unlike Sidorov et al. (2024), we include DFNet and LENS in NRP, although they use offline NVS. Our proposed method outperforms DSAC* and ACE by exploiting the additional information from the rendered novel views. Our ‘GLPoI’ beats ‘GLACE’ and achieves the state of the art.
The results of the Cambridge Landmarks datasets are shown in Table 2. Since apparent gaps exist between scene coordinate regression methods and absolute camera pose regression methods, we only list the results of SCR and NRP methods. We come to a similar conclusion as that of 7Scenes. Although we use additional rendered data, it can achieve training efficiency comparable to ACE. In order to intuitively compare the localization effects of different methods, we also used NVS to render the images corresponding to the poses estimated by different localization methods, as shown in Figure  5. The upper right is the original image, and the lower left is the image rendered using poses estimated using different methods.
@@ 4.2 Localization results
4.3 Coarse-to-fine experiments of sparse input
To further evaluate the effectiveness of our method, we do an additional experiment with sparse input as mentioned in Chapter 3.5.
implemente details: We use pre-trained MVSplatChen et al. (2024b) as the sparse NVS model. For datasets like 7Scenes, it takes thousands of images to train a small indoor scene with a scale of only several meters. To simulate the sparse input, we uniformly re-sample from the input data every 50 frames. For Scene ‘heads’, we keep only 20 frames. For outdoor datasets like Cambridge Landmarks, we split the data into multiple clusters according to the ground truth translations of camera pose (4 in the experiment) and used only one cluster for training.
The numerical results are shown in Table 3, case ‘base’ denotes the sparse input circumstances with the baseline model. Case ‘coarse’ is the self-pruning step of our method only using rendered data; we still use grid sampling as the novel pose sampling method. Case ‘c2f’ denotes the fine-tuned results of our proposed method.
Table 3: Median errors of our proposed method with sparse input on 7Scenes and Cambridge dataset.
@@ 4.3 Coarse-to-fine experiments of sparse input
4.4 Ablation of PoI
Table 4: Median errors of different implementations of PoI on 7Scenes and Cambridge dataset.
Method 7Scenes Cambridge Landmarks
base 1.1cm 0.3° 97.1% 17.7cm 0.3° 32.4%
base+poa 2.3cm 1.58° 89.6% 17.6cm 0.3° 32.2%
base+por 1.4cm 0.46° 95.9% 18.0cm 0.3° 30.4%
base+poi 1.0cm 0.3° 98.9% 16.4cm 0.3° 33.1%
To evaluate the effectiveness of our PoI approach, we conducted some experiments on PoI in different settings. As shown in Table 4, we set the training process using only query data from the training set as the case ‘base’. In this case, the training setting is similar to ACE’s. Case ‘base+poa’ indicates the training with data from the training set and all rendered pixels of the proposed novel pose rendering method. Case ‘base+por’ denotes the pixel-of-random method, which uses randomly sampled pixels from the NVS, and all else are the same with case ‘base+poi’. Case ‘base+poi’ is our final method. From the results, we may find that if we directly use sampled images and the training data without filtering, the results will be worse than the baseline (7scenes). It is easy to understand that because the mapping process is filled with low-quality pixels, it would misguide the network. If we randomly sample some pixels, we still won’t get a good estimate.
5 Conclusion
In this paper, we propose a pixel of interest filter for scene coordinate regression. The filter is designed for non-end-to-end methods, which enjoy good converging speed. With the filter, we also designed a coarse-to-fine pipeline for sparse input scenarios. We conducted experiments on indoor and outdoor datasets and achieved state-of-the-art camera pose estimation with comparable training time.

### v3

@@ III-D Filtering Strategy in PoI
The filter will remove novel views features of outlier prediction. The remaining features are FoI, and the corresponding pixels of FoI are PoI. Figure 3 shows an example of the PoI results in the 7Scenes and Cambridge Landmarks datasets. After filtering, we combine and shuffle the features $F$ and FoI and put them into the scene-specific MLP Head to estimate the scene coordinates. Each pixel is saved with its corresponding frame ID and pixel coordinates to ensure realignment after shuffling.
@@
It is worth mentioning that we have set a dynamic weight for the loss of rendered pixels. Because in the early stage of training, we want the model to converge quickly. After determining the PoI, we gradually reduce the weight of the loss of PoI from 1 to 0.01, while for the pixels from query images, we set the weight to 1 during the entire training process. The loss function is designed as follows:
@@
where $\mathcal{L}_{rep}$ means the reprojection loss, T denotes query data, $\tilde{\omega}$ denotes the dynamic weight of PoI loss, changing from $\omega_{max}$ (set to 1) to $\omega_{min}$ (set to 0.01). $I_{iter}$ denotes the current iteration number and $N_{iter}$ is the total iterations. All rendered pixels are initially set as PoI. As the training progresses, we rule out outlier prediction points from PoI. At the end of the training, the choice of PoI and the loss weight of PoI are fixed.
Figure 3: An example of the results of PoI method in dataset 7Scenes and Cambridge Landmarks. To highlight the determined pixels of interest, we scale up the ‘Value’ (V) of the HSV representation of the images.
@@
III-E Coarse-to-Fine Strategy for Extreme Sparse Input
Sparse-input visual localization (as few as 4-5 frames for some indoor scenes) presents significant challenges, as CPR and SCR techniques exhibit limitations in estimating unobserved scene regions. Their training paradigm is mainly based on 2D RGB input and lacks geometric constraints. To address this fundamental limitation, we propose leveraging sparse-input NVS to strategically extend the training views.
The primary technical challenge resides in the compromised visual quality of sparse-view NVS outputs compared to their dense-view counterparts. Direct application of the original Point of Interest (PoI) framework would disproportionately weight the implicit neural mapping process toward synthetic renderings, potentially compromising localization accuracy. Our solution introduces a novel coarse-to-fine training paradigm (illustrated in Figure 4) that systematically integrates synthetic and real data.
The initial coarse phase adopts the PoI module while exclusively utilizing the rendered images. A critical innovation in this stage is the implementation of a self-pruning mechanism with an elevated error tolerance threshold for reprojection loss filtering. The subsequent refinement stage employs curriculum learning to integrate query data with high-confidence synthetic samples progressively. Notably, we strategically maintain a reduced learning rate during fine-tuning compared with the coarse stage to enable detail-oriented optimization. This dual-phase approach eliminates filtering requirements in final training iterations, permitting full data utilization.
Refer to caption Refer to caption
(a) Coarse stage results. (b) Fine stage results.
Figure 5: The localization trajectories of the coarse-to-fine method for sparse view circumstances.
We construct the mesh based on the estimated scene coordinates of the coarse stage and the fine stage and visualize the camera pose estimation results in Figure 5. We may find that in the coarse stage, the pose estimation error is relatively large, and the quality of the reconstructed details is low. In the refined stage, the performance is much better.
Our framework demonstrates robust performance through comprehensive experimental validation across both indoor and outdoor benchmarking datasets. Quantitative evaluations and implementation specifics, including ablation studies on threshold selection and learning rate scheduling, are detailed in Section IV-D.
IV Experiment
TABLE I: Median errors of camera pose regression methods and scene coordinate regression methods on the 7Scenes dataset [16]. We bold the best results.
@@ IV Experiment
IV-A Experiments Setup
TABLE II: Results on Cambridge Landmarks, because of the obvious gap between SCR-based methods and CPR-based methods, we only list SCR-based methods. column ‘Mapping time’ shows the training time of these methods (the GS model can be trained offline, and the training time of the GS model is shorter than PoI method), and column ‘Mapping size’ is the memory consumption for saving the parameters of the network. We bold the best result for group ‘SCR’ and group ‘NRP’ separately.
@@ IV-A Experiments Setup
SCR SANet [38] Yes  1min  260M 32/0.5 32/0.5 10/0.5 16/0.6 328/2 84/0.8
SRC [39] Yes 2min 40M 39/0.7 38/0.5 19/1 31/1.0 81/0.5 42/0.7
DSAC* [13] No 15h 28M 18/0.3 21/0.4 5/0.3 15/0.6 34/0.2 19/0.4
Poker [18] No 20min 16M 18/0.3 25/0.5 5/0.3 9/0.3 28/0.1 17/0.3
EGFS [33] No 21min 9M 14/0.3 28/0.1 19/0.4 5/0.2 10/0.3 15/0.3
GLACE [34] No 20min 13M 19/0.3 17/0.4 4/0.2 9/0.3 19/0.1 14/0.3
NRP LENS [11] No - - 33/0.5 44/0.9 27/1.6 53/1.6 - -
GSplatLoc [40] No - - 27/0.46 20/0.71 5/0.36 16/0.61 - -
@@
The proposed framework is evaluated on two publicly available benchmarks: the Microsoft 7Scenes dataset [16] (with SFM-derived pseudo ground truth) and the Cambridge Landmarks dataset [20]. The methodology operates exclusively on RGB inputs paired with camera extrinsic parameters, excluding depth information from 7Scenes and structural reconstruction priors from Cambridge Landmarks. Full-resolution RGB images are preserved to ensure precise pose estimation capabilities.
To optimize computational efficiency, pose sampling and novel view synthesis are preprocessed offline, with generated camera poses and rendered images stored on disk. These precomputed rendered images are loaded with the original training data during the training phase. A specialized strategy is implemented for the ‘Kitchen’ scene, where training samples are partitioned into two clusters based on camera pose distributions. Following the ACE-inspired protocol, two distinct models are trained on these clusters. At inference time, the final pose estimation is selected from the model demonstrating superior RANSAC consensus through a higher proportion of inlier correspondences. The filter parameters $\tau_{g}$ and $\tau_{r}$ were set to 0.0005 and 5, respectively, for all scenes. All experiments utilized NVIDIA V100 GPUs. PoI training employed a single GPU, whereas GLPoI training used four GPUs in a distributed data-parallel setup to expedite convergence.
IV-B Quantitative Results
The comparison of median translation and rotation errors between our proposed methods with different camera pose regression methods (top), scene coordinate regression methods (middle), and neural render pose estimation methods (bottom) in dataset 7Scenes is shown in Table I. Generally speaking, scene coordinate regression methods outperform absolute camera pose regression methods in both translations and orientations. Unlike [40], we include DFNet and LENS in NRP, although they use offline NVS. Our proposed method outperforms DSAC* and ACE by exploiting the additional information from the rendered novel views. Our ‘GLPoI’ beats ‘GLACE’ and achieves the state of the art.
The results of the Cambridge Landmarks datasets are shown in Table II. Since apparent gaps exist between scene coordinate regression methods and absolute camera pose regression methods, we only list the results of SCR and NRP methods. We come to a similar conclusion as that of 7Scenes. Although we use additional rendered data, it can achieve training efficiency comparable to ACE. To intuitively compare the localization effects of different methods, we also used NVS to render the images corresponding to the poses estimated by different localization methods, as shown in Figure 6. The upper right is the original image, and the lower left is the image rendered using poses estimated using different methods.
@@ IV-B Quantitative Results
Figure 7: Translation error of PoI and C2F with different number of training frames for scence ‘7scenes:fire’.
IV-C Ablation of PoI
To evaluate the effectiveness of our PoI approach, we conducted some experiments in different settings. As shown in Table III, we set the training process using only query data from the training set as the case ‘base’. In this case, the training setting is similar to ACE’s. Case ‘base+poa’ indicates the training with data from the training set and all rendered pixels of the proposed novel pose rendering method. Case ‘base+por’ denotes the pixel-of-random method, which uses randomly sampled pixels from the NVS, and all else are the same as case ‘base+poi’. Case ‘base+poi’ is our final method. The results may show that if we directly use sampled images and the training data without filtering, the results will be worse than the baseline (7scenes). It is easy to understand that because the mapping process is filled with low-quality pixels, which would misguide the network. A concern has been raised that enhanced performance in Novel View Synthesis (NVS) might inversely correlate with the effectiveness of PoI metrics, as improved rendering quality could theoretically reduce the necessity of PoI metrics. We also conduct experiments with different NVS baselines. Case ‘nerf+poi’ represents the implementation of NeRF as the NVS methodology. Experimental results demonstrate that employing 3D Gaussian Splatting (3DGS) for NVS yields superior outcomes compared to NeRF. This observation suggests that the improvement of the effectiveness of the PoI metric will increase with the enhancement of the performance of the NVS method. Experiments also show that images rendered by 3DGS, like those rendered by NeRF-based methods, cannot meet the requirements for direct use by SCR.
IV-D Coarse-to-Fine Experiments with Extremely Sparse Input
To further evaluate the effectiveness of our method, we conduct an additional experiment with sparse input as mentioned in Section III-E. The comparison of translation error between PoI and C2F for scene ‘fire’ is shown in Figure 7. We can see that C2F has an advantage when the input data is sparse.
We use pre-trained MVSplat[41] as the sparse NVS model. To simulate the sparse input, we uniformly resample from the input data every 200 frames. For outdoor datasets like Cambridge Landmarks, we split the data into multiple clusters according to the ground truth translations of camera pose (4 in the experiment) and used only one cluster for training.
The numerical results are shown in Table IV, case ‘base’ denotes the sparse input circumstances with the baseline model. Case ‘coarse’ is the self-pruning step of our method only using rendered data; we still use grid sampling as the novel pose sampling method. Case ‘c2f’ denotes the fine-tuned results of our proposed method.
TABLE III: Median errors of different implementations of PoI method on 7Scenes and Cambridge dataset.
Method 7Scenes Cambridge Landmarks
base 1.1cm 0.3° 97.1% 17.7cm 0.3° 32.4%
base+poa 2.3cm 1.58° 89.6% 17.6cm 0.3° 32.2%
base+por 1.4cm 0.46° 95.9% 18.0cm 0.3° 30.4%
nerf+poi 1.1cm 0.4° 95.3% 17.2cm 0.3° 31.6%
base+poi 1.0cm 0.3° 98.9% 16.4cm 0.3° 33.1%
TABLE IV: Median errors of our proposed method with sparse input on 7Scenes and Cambridge dataset.
@@ IV-D Coarse-to-Fine Experiments with Extremely Sparse Input
V Conclusion
In this paper, we propose a PoI method for SCR-based visual localization methods. PoI applies NVS methods, including NeRF and 3DGS, to render synthetic views for training. Experiments show that the localization accuracy improves with the improvement of NVS quality. Furthermore, we designed a coarse-to-fine pipeline for extremely sparse input scenarios. We conducted experiments on indoor and outdoor datasets and achieved state-of-the-art camera pose estimation with comparable training time.